OpenStudy (anonymous):

For f(x) = sin^2 x - cos x and 0≤x≤π, find the value(s) of x for which.... 1. f(x) has a local maximum. Write EXACT answer. If none, write NA. x=_______ 2. f(x) has a local minimum. Write EXACT answers in ascending order. x=_______, x=_______ 3. f(x) has a global maximum. If there are none, write NA. Write EXACT answer. x=_______ 4. f(x) has a global minimum. If there are none, write NA. Write EXACT answer. x=________ so find derivative first? then set equal to 0? :/ thanks!!!

OpenStudy (anonymous):

not quite sure how to find the derivative of this one :/

ganeshie8 (ganeshie8):

use chain rule

ganeshie8 (ganeshie8):

\(\large f(x) = \sin^2 x - \cos x \)

ganeshie8 (ganeshie8):

u familiar wid chain rule right ?

OpenStudy (anonymous):

yes:) so would it be sinx * 2cosx + 1 ?

OpenStudy (anonymous):

not sure if i did that right though :/

ganeshie8 (ganeshie8):

yup ! whats wid +1 in the end ?

ganeshie8 (ganeshie8):

Oh, you tried to factor out sinx is it ?

ganeshie8 (ganeshie8):

you're right :)

OpenStudy (anonymous):

yay!! :D

ganeshie8 (ganeshie8):

what will be the next step ?

OpenStudy (anonymous):

umm set equal to 0? :/

ganeshie8 (ganeshie8):

yes!

ganeshie8 (ganeshie8):

\(\large f(x) = \sin^2 x - \cos x \) \(\large f'(x) = 2\sin x(\sin x)' + \sin x \) \(\large \qquad= 2\sin x \cos x + \sin x \) \(\large \qquad= \sin x (2\cos x + 1) \)

ganeshie8 (ganeshie8):

corrected the typo^

OpenStudy (anonymous):

so we get this? \[sinx(2cosx+1) =0\] and you get sinx =0 and 2cos x + 1 =0 ?

OpenStudy (kainui):

Looks like you have 3 points to solve for. Go for it!

ganeshie8 (ganeshie8):

Yes ! keep going :)

OpenStudy (anonymous):

three points? :/ and so x=0 for sinx =0 and cos x = 1/2... x= 1/2cos ?

ganeshie8 (ganeshie8):



ganeshie8 (ganeshie8):

sin(x) = height of point on unit circle

ganeshie8 (ganeshie8):

\(\large \sin x = 0\) when, \(x = 0\) and \(x = \pi\), right ?

OpenStudy (anonymous):

yes:)

ganeshie8 (ganeshie8):

so that gives u two solutions work out the other factor : \(\large 2\cos x + 1 = 0 \) \(\large \cos x = -\frac{1}{2}\) \(\large x = ?\)

OpenStudy (anonymous):

\[-\frac{ 1 }{ 2\cos }\] ?

ganeshie8 (ganeshie8):

lol no

ganeshie8 (ganeshie8):

\(\large \cos x\) is a SINGLE FUNCTION. you cannot break it like that.

ganeshie8 (ganeshie8):

for example : if \(\large f(x) = \frac{1}{2}\) then, do u say : \(\large x = \frac{1}{2f}\) ?

OpenStudy (anonymous):

ohh okay ahh i see so would it be 2* cos?

ganeshie8 (ganeshie8):

\(\large \cos x = -\frac{1}{2} \) \(\large x = \cos^{-1}(-\frac{1}{2})\)

ganeshie8 (ganeshie8):

\(\large \cos^{-1}\) is the inverse function of \(\large \cos\)

OpenStudy (anonymous):

ohh okay:)

ganeshie8 (ganeshie8):

that gives \(\large x = \frac{2\pi}{3}\)

ganeshie8 (ganeshie8):